OpenStudy (anonymous):

Please help. Solve the equation (4x+1)(x-3)=9

OpenStudy (tkhunny):

Write it in standard form. This requires "= 0". 9 is no good in that spot.

OpenStudy (mertsj):

1. Multiply the two binomials 2. Subtract 9 from both sides. 3. Try to factor 4. Set each factor equal to 0 5. Solve each equation.

OpenStudy (mertsj):

If it won't factor, plug it into the quadratic formula.

OpenStudy (anonymous):

FOIL the equation to get this; \[4X^2+X-12X-3=9\] So we get \[4X^2-11X=6\] Bring 6 to the left side and set it equal to 0 \[4X^2-11X-6=0\] Now factor to get \[(2x-6)(2X+1)=0\] Set each equal to 0 and solve \[2X-6=0\] \[2X=6\] \[X=3\] \[2X+1=0\] \[2X=-1\] \[X=-(1/2)\] So \[X= -(1/2),3\]

OpenStudy (mertsj):

@AEYES13 -3-9=-12

OpenStudy (anonymous):

@Mertsj good catch it would have worked had that actually worked at it would be positive 12 at the top

OpenStudy (mertsj):

Yes and will not factor.